OpenStudy (anonymous):

If f(x) = 6 – 5x and g(x) = 4x – 1, evaluate f(x) – g(x) for x = –2.

OpenStudy (anonymous):

h(x) =f(x) - g(x)=6 – 5x -(4x – 1)= 6- 5x -4 x +1=-9 x +7

OpenStudy (anonymous):

calculate h(-2)

OpenStudy (anonymous):

h(-2) = -9(-2) +7 = 18 + 7 =25